A company plans to interview 10 recent graduates for possible employment. The company has four positions open. How many groups of four can the company select?


1
Expert's answer
2021-11-30T09:53:13-0500

"N = \frac{10!}{4!(10-4)!} = \frac{7\times 8 \times 9 \times 10}{2 \times 3 \times 4} = 210"